OpenStudy (anonymous):

Predict the next two numbers in this sequence. 1/4, 1/9, 1/16, 1/25,... A. 1/36, 1/49 B. 1/36, 1/47 C. 1/50, 1/100 D. 1/50, 1/75

OpenStudy (zehanz):

1/4 = 1/2 * 1/2 1/9 = 1/3 * 1/3 1/16 = 1/4 * 1/4 1/25 = 1/5 * 1/5 . .